Decide how many solutions this equation has: x^2-4x+4=0
It's my summer assignment for a Pre AP class & it's a major test grade. Thank you so much for your help!!

Answer by checkley75(3666)

X^2-4X+4=0
(X-2)(X-2)=0
X-2=0
X=2 ANSWER. THERE IS ONLY ONE ANSWER (2).